radomes.openskiesproject.org
cierrashepherd2003
cierrashepherd2003 cierrashepherd2003
  • 20-02-2016
  • Mathematics
Answered

If P(blue) is 0.3, what is P(not blue)?\

A0.3
B0.5
C0.6
D0.7

Verified Answer :

olemakpadu
olemakpadu olemakpadu
  • 20-02-2016

Probability of Event Occurring + Probability of Event not occurring = 1.

Therefore:

P(blue) + P(not blue) = 1.

Therefore P(not blue) = 1 - P(blue)

P(blue) = 0.3

P(not blue) = 1 - P(blue) =1 -  0.3 = 0.7

P(not blue) = 0.7

Option D.

Answer Link

Related Questions

How many solutions does the equation y 12 y 10 2 have
how do t cells differ from b cells
You want to decorate the bottom of an aquarium with brightly colored rocks of different shapes and sizes this describes a
What is 225 written are a fraction in simplest form
A triangle has two sides that measure 10 in and 8 in which could be the measure of the third side a 20 in b 17 in c 2 in d 1 in
Whom is Arete a descendent of
A car leaves a town traveling at 60 miles per hour Two hours later a second car leaves the town traveling the same route and overtakes the first car in 6 hours
pilgrims faced many hardships on their voyage What were three hardshpis they faced aboard the Mayflower
what is the measure of an exterior angle of a regular 7sided polygon
23x1 polynomial operations